OpenStudy (anonymous):

Ln 1=x I think it's 0 but now 100% sure Please help

OpenStudy (anonymous):

It's zero. How do you get rid of a natural log? You put an 'e' underneath it. e^ln(1) = e^x. 1 = e^x. Any positive number to what power equals 1? Any positive number to the power of 0 equals 1. 3^0 = 1, 100^0=1, 5002346^0 = 1. e^0 = 1.

OpenStudy (anonymous):

Thank you I though it was 0 or 1 but wasn't sure